What is the product of the complex numbers (3+i) and (3-i)?
A) 8
B) 10
C) 9-i
D) 10-6i

Found 2 solutions by Edwin McCravy, stanbon:
Answer by Edwin McCravy(20086)

(3 + i)(3 - i) = 9 - 3i + 3i - i² = 9 - i² = 9 - (-1)² = 9 - 1 = 8

Answer by stanbon(75887)

What is the product of the complex numbers
(3+i) and (3-i)
---
Use FOIL:
F: 3*3 = 9
O: 3*-i = -3i
I; 3*i = 3i
L: i*-i = i^4 = 1
----
Combine to get 9+1 = 10
=============================
Cheers,
Stan H.
=============================
A) 8
B) 10
C) 9-i
D) 10-6i